### v3

# Verifiable End-to-End Delegated Variational Quantum Algorithms

Abstract
Variational quantum algorithms (VQAs) have emerged as promising candidates for solving complex optimization and machine learning tasks on near-term quantum hardware. However, executing quantum operations remains challenging for small-scale users because of several hardware constraints, making it desirable to delegate parts of the computation to more powerful quantum devices. In this work, we introduce a framework for delegated variational quantum algorithms (DVQAs), where a client with limited quantum capabilities delegates the execution of a VQA to a more powerful quantum server. In particular, we introduce a protocol that enables a client to delegate a variational quantum algorithm to a server while ensuring that the input, the output and also the computation itself remain secret. Additionally, if the protocol does not abort, the client can be certain that the computation outcome is indeed correct. This work builds on the general verification protocol introduced by Fitzimons and Kashefi (2017), tailoring it to VQAs. Our approach first proposes a verifiable protocol for delegating the quantum computation at each optimization step of a VQA, and then combines the iterative steps into an error-resilient optimization process that offers end-to-end verifiable algorithm execution. We also simulate the performance of our protocol tackling the Transverse Field Ising Model. Our results demonstrate that secure delegation of variational quantum algorithms is a realistic solution for near-term quantum networks, paving the way for practical quantum cloud computing applications.
preprint: APS/123-QED
I Introduction
Quantum computing is poised to revolutionize information processing by leveraging the principles of quantum mechanics to solve problems that are intractable for classical computers. With advancements in quantum hardware, software, and algorithms, we are witnessing rapid progress toward practical quantum computation. However, the realization of fault-tolerant, large-scale quantum computers remains a long-term goal. In the near term, noisy intermediate-scale quantum (NISQ) devices—quantum processors with tens to hundreds of qubits but without robust error correction—represent the frontier of quantum technology. These NISQ devices offer promising opportunities for obtaining a quantum advantage in specific computational tasks, such as quantum simulation, optimization, and machine learning. Yet, their inherent noise and limited coherence times present significant challenges.
Variational quantum algorithms (VQA) have emerged as a promising approach to leverage NISQ devices to solve complex computational problems [5]. They are hybrid quantum-classical algorithms, in which a quantum processor executes a circuit with adjustable parameters, and a classical optimizer iteratively updates these parameters to minimize a given cost function. By combining quantum circuits with classical optimization techniques, VQAs can efficiently explore high-dimensional solution spaces while mitigating the limitations of NISQ hardware.
As quantum processors become more accessible via cloud-based platforms, it is natural to consider the concept of delegated variational quantum algorithms (DVQAs). Such a process is envisioned to enable a classical client to delegate complex quantum computational tasks to a remote quantum server while maintaining control over the optimization process. Ensuring the privacy and correctness of delegated computations is crucial, particularly in scenarios where the client lacks direct access to the quantum hardware.
Most delegated quantum computation protocols are instantiated in the measurement-based quantum computing (MBQC) model, which naturally allows for blindness and verifiability of the computation [4, 2, 10, 11]. Compared to the circuit model, MBQC inherently supports these features due to the one-way structure and the decoupling of resource state preparation from the actual (measurement-based) computation. Moreover, MBQC is particularly well suited to photonic implementations and quantum network architectures, further reinforcing its advantages for delegated quantum computation.
In the context of variational quantum algorithms, a direct application of the aforementioned established delegation methods is not feasible. Due to the hybrid nature of VQAs, which involves classical optimization, a tailored framework for delegating such algorithms is required. This article explores the potential of measurement-based delegated variational quantum algorithms (MB-DVQAs). While previous studies have already provided the first investigations of MB-DVQAs [15, 27, 30, 31], they focused on the shot-level verifiability of single iterations of the quantum variational procedure and therefore did not achieve verifiability for the entire optimization process.
We give an alternative protocol to achieve individual optimization step verifiability, and then propose a verifiable end-to-end MB-DVQA, encompassing both quantum and classical computations involved.
Overview of the contributions
  1. 1.
    We show how to verify one step of an MB-DVQA (Protocol III.1) by using a trapification scheme that alternates computation and test rounds [13], and prove that it succeeds with probability that grows exponentially with the number of total rounds (Theorem III.2).
  2. 2.
    We propose a new MB-DVQA using a custom gradient-descent optimizer (Protocol III.2) that reruns highly attacked iterations without aborting the full optimization and show that it is verifiable (Theorem III.3).
  3. 3.
    We benchmark the custom optimizer outlined in Protocol III.2 through numerical simulations of a Variational Quantum Eigensolver (VQE) [19] that finds the ground state energy of the Transverse Field Ising Hamiltonian, incorporating simulated gradient perturbations that mimic adversarial attacks.
II Background
A central component in VQAs is a Parametrized Quantum Circuit (PQC). This circuit (also called Ansatz), is associated with a unitary $U(\theta)$, where $\theta$ is a vector of parametrized angles of dimension $N_{P}$. The main goal of a VQA is to minimize a cost function $f$ represented as:
$f(\theta)=Tr[OU(\theta)\rho_{in}U^{\dagger}(\theta)]$
where $\rho_{in}$ is the initial state and $O$ is the observable describing the problem to solve. In particular, $O$ is decomposed as sum of $N_{o}$ local Pauli operators $P_{i}$:
$O=\sum_{i}^{N_{o}}O_{i}=\sum_{i}^{N_{o}}c_{i}P_{i}$
We will focus on gradient-based optimization, making use of the Parameter Shift Rule [18, 25] to compute an approximation $g$ of the gradient $\nabla f$. For the resulting vector $g=(g_{1}(\theta),\dots,g_{N_{p}}(\theta))$, it holds that for every element of the gradient vector the partial derivative is $\frac{\partial f}{\partial\theta}=\frac{1}{2}[f_{+}-f_{-}]$, where:
$f_{+}=f(\theta+\frac{\pi}{2})\text{ and }f_{-}=f(\theta-\frac{\pi}{2}).$
This means that for a given PQC with $N_{P}$ parameters, the total number of function evaluations per step is $2N_{P}$. Using a fixed amount of measurement shots per function evaluation $N_{s}$, we require a total number of $2N_{P}N_{s}$ shots to perform one step of the optimization process.
Till recently, all proposals of VQEs were based on the circuit model of computation. In the last years, variational quantum algorithms based on measurement-based quantum computation have also been proposed [7, 21, 24, 16, 28, 17, 22]. MBQC [23] differs significantly from the conventional gate-based approach; instead of executing quantum gates sequentially, MBQC begins with the preparation of a highly entangled state, and the computation is then carried out through a sequence of adaptive quantum measurements that depend on the outcomes of the previous measurements.
[…]
In a previous study of delegated bounded-error quantum polynomial computation [13], the proposed verification scheme interleaves $d$ computation rounds and $t$ test rounds (for a total of $n=d+t$ rounds), rather than inserting traps into the graph $G$ associated with the desired computation. The latter approach results in an enlarged graph, whereas interleaving rounds maintain the original graph size. In the computation rounds, the Client prepares and sends the qubits and the measurements associated to the desired computation unitary $U$ (defined by graph $G$), while the test rounds instead are just made up of traps and dummy qubits. In particular, [13] makes use of a $c$-coloring of $G$ to place traps and dummy qubits. By having the same graph structure for both computation and test rounds, the Server cannot distinguish them due to blindness inherited by previous works [4].
III Methods
Our goal is to provide an MB-DVQA that converges and is at the same time blind, secure and verifiable. Since blindness and security are directly inherited from previous works on delegated MBQC [6], the only thing that remains to address is verifiability. We first propose a verifiable protocol (Protocol III.1) computing the gradients of one MB-DVQA step using the parameter shift rule. In particular, we provide a bound (Lemma III.1) on the gradient’s relative error in case of adversarial behavior, namely when cost function evaluations are affected by a certain number of corrupted computation rounds. Subsequently, we make use of this result to come up with a trapification scheme based on [13], yielding an undetected highly corrupted outcome with a very low probability (Theorem III.2). We then exploit this protocol within the context of a complete MB-DVQA using gradient descent to minimize the cost function. Specifically, we modify the classical optimization part of the VQA by re-performing highly attacked steps detected by the step-level protocol and prove convergence and verifiability (Theorem III.3) for the full MB-DVQA protocol (Protocol III.2).
III.1 Verifiability for one step of the MB-DVQA
The aim of Protocol III.1 is to delegate the quantum computation required to compute the gradient of the cost function with a relative error bounded by a target value $e_{th}$. This quantum computation consists of all the $2N_{P}$ evaluations of the cost function $f$, each performed $N_{s}$ times.
To provide verifiability for this protocol and allow the client to check for the correctness of the computation, we use a trapification scheme proposed in [13]. In that work, majority vote was used to bound the probability of failure of the verification process. Here instead, due to the sampling nature of PQCs, we consider the computation outcome through the expectation value of observables, which are typically used to evaluate the cost function. We therefore execute $d=2N_{P}N_{s}$ computation rounds and $t$ test rounds shared across different quantum computations (i.e. distinct $f$ evaluations). This allows us to reduce the amount of required test rounds compared to the case where each quantum computation is delegated separately.
{algorithm}
Verifiable Step of the MB-DVQA
1:Inputs: $G$ and associated $c$-coloring, $\theta=\{\theta_{i}\}_{i=1}^{N_{P}}$, $N_{s}$, $e_{th}$.
2:Protocol:
  1. 1.
    Suppose $n=d+t$ the number of rounds, where $d=2N_{P}N_{s}$ are the computation and $t$ are the test rounds and the Client chooses at random which round belongs in each set. For each round:
    • if it is a test round, the client chooses a random graph coloring and sends the trap and dummy qubits to the server.
    • if it is a computation round, the client sends the qubits associated to $G$.
    • the server performs the blind measurement based computation and returns the output.
  2. 2.
    The client gathers and decodes the server’s output:
    • in case she detects a trap failure, she aborts.
    • otherwise, she accepts, obtaining $g(\theta)$ with a relative error up to $e_{th}$.
In order to prove verifiability for our protocol, we first establish a connection between the relative error in the gradient approximation and the adversarial attacks, represented in our framework by means of the number of corrupted computation rounds $\delta$.
Lemma III.1.
[…]
$e\leq\frac{\sum_{i}|c_{i}|}{\epsilon_{0}N_{s}}\delta$ (1)
Proof.
When we perform one execution of the PQC, the state collapses to an eigenstate of the measured observable. Therefore, the cost function can be expressed as:
$f=\frac{N_{o}}{N_{s}}\sum_{i=1}^{N_{\text{o}}}c_{i}\sum_{k=1}^{N_{s}/N_{o}}\lambda^{i}_{k}$
Let’s consider a single eigenvalue $\lambda^{i}_{k}$ of the local Pauli $P_{i}$ measured at the $k$-th round. Due to attacks, in the worst-case scenario, we instead measure a corrupted eigenvalue equal to the most distant w.r.t. the correct $\lambda^{i}_{k}$. One can bound the difference between the correct and the worst-case corrupted eigenvalues as follows:
$\lambda^{i}_{k}-\hat{\lambda}^{i}_{k}=\lambda^{i}_{k}-\max_{j}|{\lambda^{i}_{k}-\lambda^{i}_{j}}|\leq|{\lambda^{i}_{max}-\lambda^{i}_{min}}|\leq 2$
In the presence of $\delta^{\prime}$ corrupted computation rounds per function evaluation, we assume that these rounds are equally distributed among the different local Pauli expectation value estimations, resulting in $\delta^{\prime}/N_{o}$ corrupted computation rounds for each estimation. Therefore, the absolute difference between the ideal function $f$ and its corrupted version can be bounded as follows:
$|f-\hat{f}|=\frac{N_{o}}{N_{s}}\sum_{i=1}^{N_{\text{o}}}|c_{i}\sum_{k=1}^{N_{s}/N_{o}}(\lambda^{i}_{k}-\hat{\lambda}^{i}_{k})|\leq\frac{2\delta^{\prime}}{N_{s}}\sum_{i=1}^{N_{\text{o}}}|c_{i}|$
The relative error in the gradient computation obtained with the parameter shift rule will be:
$e$ $=\frac{\|g(\theta)-\hat{g}(\theta)\|}{\|g(\theta)\|}=\frac{1}{\|g(\theta)\|}\sum_{i=0}^{N_{P}}|g_{i}(\theta)-\hat{g}_{i}(\theta)|$
$=\frac{1}{2\|g(\theta)\|}\sum_{i=0}^{N_{P}}|(f_{+}-\hat{f}_{+})-(f_{-}-\hat{f}_{-})|$
The total number corrupted computation rounds per step, distributed across the function evaluations required for all the components of the gradient vector, is expressed as $\delta=\sum_{i=0}^{N_{P}}\delta_{i},\quad\text{where }\delta_{i}=\delta_{i+}+\delta_{i-}$. Therefore, the absolute error can be bounded as follows:
$\|g(\theta)-\hat{g}(\theta)\|$ $\leq\frac{\sum_{i}|c_{i}|}{N_{s}}\sum_{i=0}^{N_{P}}|\delta_{i+}+\delta_{i-}|=\frac{\sum_{i}|c_{i}|}{N_{s}}\delta$
Since we cannot access the real value $\|g\|$, we can assume a lower bound $\epsilon_{0}$ leading to the Lemma’s result. ∎
Given a desired target relative error $e_{th}$, Lemma III.1 provides a bound on the tolerated corrupted computation rounds $\delta_{max}$ affecting one Step:
$\delta\leq\delta_{max}=\frac{e_{th}\epsilon_{o}N_{s}}{\sum_{i}|c_{i}|}=\frac{e_{th}\epsilon_{0}d}{2N_{P}\sum_{i}|c_{i}|}=wd$ (2)
where $w=\frac{e_{th}\epsilon_{0}}{2N_{P}\sum_{i}|c_{i}|}$.
Now that we have a bound on $\delta$, we can proceed with showing verifiability of a MB-DVQA step.
Theorem III.2.
An MB-DVQA step following Protocol III.1 computes the gradient of $f$ with relative error up to $e_{th}$ and probability of failure decreasing exponentially with the number of rounds $n$.
Proof.
Given a total of $m$ attacked rounds, let $X_{d}\approx\text{Binomial}(m,\frac{d}{n})$ be the random variable describing the probability of computation rounds being attacked. Also, let $X_{t}\approx\text{Binomial}\left(\left(\frac{m}{n}-\epsilon_{1}\right)t,\frac{1}{c}\right)$ be the random variable related to the number of attacks detected by the traps in the trapification scheme (with some $\epsilon_{1}>0$). Therefore, the probability of failure will be the probability of having more than $\delta_{max}$ corrupted computation rounds and no detected attacks by any trap:
[…]
Let’s define $X^{\prime}\approx\text{Binomial}(m,\frac{t}{n})$ to be the probability distribution describing the attacked test rounds. By the law of total probability, the probability to have no detected attacks in the test rounds is:
$\text{Pr}(X_{t}<1)$ $=\text{Pr}\left[X_{t}=0|X^{\prime}\leq\frac{m}{n}t\right]\text{Pr}\left[X^{\prime}\leq\frac{m}{n}t\right]$
$+\text{Pr}\left[X_{t}=0|X^{\prime}>\frac{m}{n}t\right]\text{Pr}\left[X^{\prime}>\frac{m}{n}t\right]$
$\leq\text{Pr}\left[X^{\prime}\leq\frac{m}{n}t\right]+\text{Pr}\left[X_{t}=0|X^{\prime}>\frac{m}{n}t\right]$
Using Hoeffding’s inequality and the fact that $(1-x)^{k}\leq e^{-xk}$, we can rewrite the above probability (with $\tau=t/n$):
$\text{Pr}(X_{t}<1)$ $\leq\text{exp}\left(\frac{-2\epsilon_{1}^{2}\tau^{2}n^{2}}{m}\right)+\text{exp}\left(-\left(\frac{m}{n}-\epsilon_{1}\right)\frac{\tau n}{c}\right)$
To bound the probability of failure we analyze it in two complementary scenarios.
A. When $m>nw$, considering that computation and test rounds are chosen at random, the number of affected computation rounds $m\frac{d}{n}$ is higher than $\delta_{max}$, leading to $\text{Pr}[X_{d}\geq\delta_{max}]=1$. Therefore we focus on the second term of Eq. (3) to bound the probability of failure.
$\text{Pr}(X_{t}<1)\leq\text{exp}\left(\frac{-2\epsilon_{1}^{2}\tau^{2}n}{w}\right)+\text{exp}\left(-\frac{\tau n}{c}\left(w-\epsilon_{1}\right)\right)$
Both terms decay rapidly to zero with respect to the number of copies if $w>\epsilon_{1}$. If $m\geq nw$, $\text{Pr}[X_{d}\geq\delta_{max}]\rightarrow 1$ but $\text{Pr}[X_{t}<1]$ decays exponentially with $n$.
B. When $m<nw$, using Hoeffding’s inequality for the first term of Eq. (3):
[…]
$\leq\text{exp}\left(-\frac{2d^{2}}{m}\left(\frac{m}{n}-w\right)^{2}\right)$
$\left(\text{exp}\left(\frac{-2\epsilon_{1}^{2}\tau^{2}n^{2}}{m}\right)+\text{exp}\left(-\left(\frac{m}{n}-\epsilon_{1}\right)\frac{\tau n}{c}\right)\right)$
Here $\forall m>0$ with $1/n>\epsilon_{1}$, $\text{Pr}[X_{t}<1]$ decreases exponentially with $n$. When there are no attacks, $m=0$ then $\text{Pr}[X_{d}\geq\delta_{max}]\rightarrow 0$, therefore we have a bound on the probability of failure decaying exponentially w.r.t. the number of copies $n$ and the number of test rounds $t$, since $d$ is fixed.
So far, the verifiability of one step of the MB-DVQA has been shown. However, it might not be clear how to choose the threshold on the tolerated gradient relative error $e_{th}$. In the next section, we will show how to appropriately set this bound to ensure both convergence of the optimization and verification of the full MB-DVQA.
III.2 Verifiable MB-DVQA with Gradient Descent
In this section we explore what happens when we build a verifiable MB-DVQA protocol (Protocol III.2) using Protocol III.1. Given limited resources, namely a total number of available rounds, the MB-DVQA protocol runs Protocol III.1 for each step of the optimization. In case a step aborts, the protocol rejects the last gradient computation and re-performs the same step; otherwise it accepts the outcome (with a relative error at most $e_{th}$) and updates the variational parameters.
In the scenario where a delegated VQA is executed with verifiability provided only at the step level, there is no guarantee regarding the final outcome of the variational optimization, namely the final value of $f$. Although highly corrupted steps can be rerun, once the MB-DVQA is terminated, we cannot clearly state if the outcome matches the optimum of the cost function $f^{*}$. Therefore, such a protocol would have no practical means of knowing whether an execution is successful. To address this gap, we introduce protocol III.2 that achieves verifiability by means of geometric convergence of the optimization procedure, as shown in Theorem III.3. This is done by properly tuning $e_{th}$ depending on the problem, a task made possible by the result achieved in the previous sections with Theorem III.2 and Lemma III.1. Given convergence guarantees, once the protocol terminates, the Client either obtains the desired outcome $f^{*}$ or receives an abort response, achieving verifiability.
{algorithm}
Verifiable MB-DVQA with gradient descent
Inputs: Initial angles $\theta^{1}$, maximum number of steps $N_{\text{iter}}$, a $c$-coloring, $N_{s}$, $e_{th}$ and learning rate $\alpha$.
Protocol:
  1. 1.
    For $k\in[1,...,N_{\text{iter}}]$, run Protocol III.1 to compute $\hat{g}(\theta^{k})$:
    • If it accepts: $\theta^{k+1}=\theta^{k}-\alpha\hat{g}(\theta^{k})$.
    • If it aborts: $\theta^{k+1}=\theta^{k}$ (i.e. re-run the same step).
  2. 2.
    Set the final angles: $\theta^{*}\leftarrow\theta^{N_{\text{iter}}}$.
  3. 3.
    If the optimization hasn’t converged, aborts; otherwise accepts returning $\hat{f}^{*}$.
Theorem III.3.
Given a $\mu$-strongly convex cost function $f$ with $L$-Lipschitz continuous gradient and a global optimum $f^{*}$, Protocol III.2 ensures verifiability, provided that the learning rate $\alpha$ and error bound on the gradient $e_{th}$ satisfy the following conditions:
$\mu\alpha(1+e_{th})[2-\alpha L(1+e_{th})]<1;\quad\alpha L<\frac{2}{1+e_{th}}$ (4)
Sketch of Proof.
As explained in detail in Appendix VI.2, a VQA using Stochastic Gradient Descent to optimize a function $f$ can converge in the presence of adversarial attacks provided that the gradient relative error remains sufficiently bounded. If no such bound exists, multiple steps with large gradient errors may occur, potentially jeopardizing the optimization and leading to divergence. To prevent this and ensure convergence, Protocol III.2 integrates Protocol III.1, which detects steps where the relative error exceeds the threshold $e_{th}$. Such steps are then re-executed until the conditions of the Theorem hold. These inequalities, derived in the full proof (Appendix VI.2), are related to standard conditions required for gradient descent optimization, namely having $10<\gamma<$ 1 such that for each optimization step $k$:
$\mathbb{E}[f^{k+1}]-f^{*}\leq\gamma(\mathbb{E}[f^{k}]-f^{*})$
Notably, when $e_{th}\rightarrow 0$, the inequalities above reduce to the standard gradient descent conditions for convergence, having $\gamma=1-\mu\alpha(2-\alpha L)$. This guarantees the average decrease of the function required for convergence and also bounds the final error neighborhood in the presence of shot-noise, albeit at the cost of an overhead of steps due to repetitions. Finally, we associate convergence guarantees with verifiability: since the Client has access to classical gradient information at each step $k$, the protocol either outputs the correct result, namely the optimal value of the cost function achieved due to convergence, or aborts. ∎
IV Simulations
We start by benchmarking the performance of Protocol III.2, to demonstrate its effectiveness. We do this by trying to find the ground state of the Transverse Field Ising Model (TFIM), a fundamental quantum spin model describing a system of interacting spins (typically on a lattice):
$H=h\sum_{i}X_{i}+\sum_{<i,j>}Z_{i}Z_{j}$
where $<i,j>$ determines the edges of the lattice and $h$ is the strength of the external transverse field. We express attacks as perturbations on the gradient vector for each step $k$ (in the context of Protocol III.2):
$\hat{g}_{i}(\theta^{k})=g_{i}(\theta^{k})+\nu_{i}^{k}$ (5)
Each $\nu_{i}^{k}$ is drawn independently from a uniform distribution $\mathcal{U}(-\Delta,\Delta)$, where $\Delta$ is a parameter tuning the impact of the perturbation. The perturbation vector is defined as:
$\nu^{k}=\begin{cases}(\nu_{1}^{k},\nu_{2}^{k},\ldots,\nu_{N_{P}}^{k}),&\text{with probability }p\\ (0,0,\ldots,0),&\text{with probability }1-p\end{cases}$
The simulation of the computation rounds in the absence of quantum hardware noise is carried out using the Qrisp programming framework [26]. For now we assume the existence of traps allowing for a negligible probability of failure and leave the simulation of the test rounds for later. In this way, we can simulate the trap detection by evaluating the relative error. Protocol III.2 re-performs heavily attacked steps without aborting the full VQE and accepts steps that fall below the error threshold.
We show examples for two-dimensional TFIMs using $p=0.7$ and an ansatz with $N_{L}$-layers and $n$ qubits, each layer $U_{l}$ is given by:
$U_{l}(\theta_{l})=\left(\prod_{i=1}^{n-1}\text{CNOT}_{i,i+1}\right)\left(\bigotimes_{i=1}^{n}R_{y}(\theta_{l}^{i})\right)$
Figure 1: Average performance over 5 runs comparing the proposed protocol and a standard VQE in presence of the described attacks with $p=0.7$. We tackle the following two-dimensional TFIM lattices: $2\times 2$ and $3\times 3$ with $h=0.2$. The dashed green line (Attack Free) refers to an ideal attack-free scenario, the yellow line (No traps) is related to a non-verifiable delegated VQE in presence of attacks and finally, the proposed protocol is shown in blue (Traps). We use $N_{s}=1000$, $N_{L}=2$ layers and a learning rate $\alpha=0.2$ for the optimization.
Figure 2: Plot of relative error $e$ versus number of trap detections $N_{TD}$ in one execution of Protocol III.1 (in logarithmic scale). Here, we choose $d=4000$ and $t=50$. $p_{attack}$ is the probability that a round is attacked and $\hat{\alpha}$ shows examples of a degree shift applied by an attacker on the measurement angles of the output qubits. The points represent average scores computed over 10 runs for each setting. As expected, when the relative error exceeds the threshold $e_{th}$ (shown with the red dashed line), the attack is detected.
From Fig. 1 we can see how a standard VQE using gradient descent diverges due to attacks, while the proposed approach, as expected, turns out to be robust and converges to the optimal solution with an overhead of steps with respect to the ideal attack-free case.
It is important to emphasize that our protocol does not depend on the aforementioned attack model. The latter has been designed to illustrate the performance of the proposed work. As reported in the previous section, Protocol III.2 is general and robust, providing tools to adjust the number of traps to ensure convergence, independent of the attack model.
To complement the numerical analysis, we provide simulations of Protocol III.1 tackling a 2-qubit TFIM. In particular, we make use of Veriphix [1], a library for verifiable blind quantum computing based on the MBQC programming framework Graphix [29]. In Fig. 2 we show how the number of test rounds $t$ can be tuned to detect highly corrupted steps depending on $e_{th}$, the threshold value required by Protocol III.2 to ensure convergence.
V Conclusion
This work offers valuable insights into the future of delegated quantum computation in the NISQ era and beyond, by proposing an end-to-end verifiable delegated variational quantum protocol. The theoretical results are further validated with simulations that tackle the Transverse Field Ising Model. Although our work is currently based on gradient descent optimization, it is readily adaptable to alternative optimization techniques. Specifically, Protocol III.1 can be modified to accommodate different optimizers, leading to appropriately revised versions of Lemma III.1 and Theorem III.2. On the other hand, with regard to the verifiability of the overall VQA, an optimizer-specific Theorem should be provided to guarantee convergence and consequentially, verifiability. Furthermore, in our protocol, we assume a fixed relative error threshold for all the steps. In future work, the threshold across the steps could be varied depending on the current gradient value and accordingly the number of traps required for each step could be adjusted, thus reducing the total number of delegated computations. Another extension of this work could involve an improvement of the trapification procedure, for instance exploiting the scheme based on stabiliser testing proposed in [12].
Finally, although circuit-based DVQA protocols have already been proposed (such as the one by Li et al. [14] based on quantum homomorphic encryption) they currently only provide blindness. It remains an open and interesting question how verification could be incorporated in such schemes, either through adaptations of our approach or via dedicated techniques like accreditation [8].
Acknowledgements.
M.I. and A.P. acknowledge support from the Federal Ministry for Economic Affairs and Climate Action (project Qompiler, grant No: 01MQ22005A), the DFG (Emmy Noether programme, grant No. 418294583) and the BMBF (project tubLANQ.0, grant No. 16KISQ087K). P.W. acknowledges support by EPSRC grant EP/T001062/1.